OpenStudy (anonymous):

how do i find the angle a line segment PQ makes with the x-axis if it is plotted in 3 dimensions? P(1,1,4) and Q(0,6,-3)

OpenStudy (anonymous):

You can use vectors since \[\cos\theta = \frac{a.b}{|a||b|}\]